OpenStudy (donnie1999):

HELP SIMILAR QUESTION! Tracy needs to find the length and width of the equipment storage room where she works. the room is 2 meters longer than it is wide. Its perimeter is 24 meters. Find the width in meters.

OpenStudy (donnie1999):

@phi

OpenStudy (phi):

do you have names for the unknowns ?

OpenStudy (donnie1999):

no

OpenStudy (phi):

can you make up some names ? what do we not know or need to find ?

OpenStudy (ethan01):

Let x= width x+2 = length 2W + 2L = P 2(x) + 2(x+2) = 24 2x + 2x + 4 = 24 4x +4 = 24 Subtract 4 from each side: 4x=20

OpenStudy (donnie1999):

l = length w= width

OpenStudy (ethan01):

Divide by 4: x= 5 meters Width x+2 = 7 meters Length Check: 2*5 + 2*7 10+14=24

OpenStudy (donnie1999):

thanks but i already figured it out :D

OpenStudy (phi):

btw, the point of doing these problems is learning how to figure out the equations, and then solving them. knowing the answer is not the point. (It's the difference between watching someone practicing football versus doing the practice...)